# Looking Beyond Visible Cues: Implicit Video Question Answering via Dual-Clue Reasoning

### v1

@@ 4 Methods
In this section, we present IRM as shown in Fig. 5, which mainly consists of the visual encoder, AIM (Action-Intent Module) (Sec. 4.1), VEM (Visual Enhanced Module) (Sec. 4.2), and the LLM.
@@ 4.1 Action Intent Module
As discussed above, based on the theories from cognitive science, we construct contextual clue information as an additional reasoning pathway to assist implicit reasoning tasks. Specifically, leveraging the powerful basic recognition and reasoning capabilities of IRM (initialized from VideoChat2 videochat2 after vision-language alignment and connection), we prompt it with instruction prompt P (shown in the left-top legend of Fig. 5P: Task: Please extract …) to generate action-intent clue candidates.
Since the visual evidence in the context is explicit, extracting these direct, shallow intentions from the actions is much less challenging than directly extracting the indirect, implicit intentions behind the question. LLM is prompted with the uniformly sampled frames of implicit visual input $v$ and the In-Context Learning prompt. Context action sets $\mathbb{A}_{c}=\{act^{1},...,act^{n}\}$ and corresponding intents set $\mathbb{I}=\{int^{1},...,int^{n}\}$ are generated, which are then embedded and concatenated as follows:
$\small X_{A}=\texttt{Cat}(f_{T}(act^{1}),...,f_{T}(act^{n})),X_{I}=\texttt{Cat}(f_{T}(int^{1}),...,f_{T}(int^{n}))$ (1)
where $f_{T}$ indicates the word embedding layer, Cat indicates concatenate. To further refine the clues, we focus on mitigating the hallucination problem of clues while reasoning out key clues beneficial for implicit reasoning. The current action clues may be influenced by hallucinations, but this can be effectively mitigated through careful verification using visual information MECD ; hallucination ; mmvu . Therefore, we incorporate cross-action-vision attention to query each action candidate with global visual information $Q_{V}$ extracted by QFormer. Specifically, the mechanism is performed by the following equation:
$\small X_{I}^{V}=\text{$\Phi_{h}^{C}$}(X_{A},f_{P}(Q_{V}))$ (2)
@@
Moreover, action-intent clue pairs contribute differently to the reasoning of the implicit question. Therefore, we introduce the causal relations classifier with implicit question $q$ to determine which intent-action pairs assist in reasoning. Specifically, we concatenate the visually enhanced $X_{I}^{V}$ and the original intent tensor $X_{I}$ with the embedded question $f_{T}(q)$ and then pass them to the causal attention classifier MECD for relationship deduction. This process is represented as:
@@
$\small\mathcal{L}_{R}=\mathcal{L}_{ce}(r,\hat{r}*r_{m})$ (4)
where $\mathcal{L}_{ce}$ indicates the cross-entropy loss, $r_{m}$ indicates the matching results. The value of $r_{m}$ is 0 when the action candidate does not match any action pairs annotation, or when the action matches while the deduced intent is incorrect according to the intent annotation. Otherwise, the value of $r_{m}$ is 1. The matching or not is based on the automatic annotations from GPT-4. During inference, the refined context action clues $X_{A}^{\prime}$ and intent clues $X_{I}^{\prime}$ can be represented as follows:
$\small X_{A}^{\prime}=\texttt{Cat}\left({X_{A}}[i]\right),X_{I}^{\prime}=\texttt{Cat}\left({X_{I}}[i]\right),~{}\hat{r}[i,0]>\hat{r}[i,1].$ (5)
@@
Figure 5: IRM Framework. IRM employs dual-stream modeling of contextual actions and intent clues as implicit reasoning chains. Dual clues are generated and refined by AIM through hallucination mitigation and clue-question relationship deduction. Meanwhile, VEM takes the refined clues to enhance visual input. Finally, LLM is prompted with dual clues to conduct implicit reasoning.
@@ 4.2 Visual Enhanced Module
As discussed above, the VEM first introduces QFormer blip2 to compress extensive visual input $f_{V}(v)$ from the visual encoder, represented as $Q_{V}$. During the compression, the instruction to Qformer is: “Extract the context clues related to the question: q”. Following the previous works videochat2 ; videollava ; videollama ; compress , compressed visual input $Q_{V}$ is then sent to visual projection layers $f_{P}$ to align with textual embeddings, represented by $X_{V}=f_{P}(Q_{V})$.
The aligned visual input is then sent to the AIM module for dual context clues reasoning. Meanwhile, the current visual extraction $X_{V}$ only implicitly suggests the need to extract key context visual clues. Therefore, for explicit visual enhancement, we establish a feedback path from AIM to VEM, utilizing context clues generated by AIM to enhance the visual extraction of VEM:
$\small X_{V}^{\prime}=\text{$\Phi_{v}^{C}$}(X_{V},\texttt{Cat}(X_{A}^{\prime},X_{I}^{\prime})).$ (6)
@@
$\small Ans=LLM(X_{A}^{\prime},X_{I}^{\prime},X_{V}^{\prime},q).$ (7)
@@ 5 Experiments
Table 1: Results on I-VQA. IRM reaches SOTA performances in both multi-choice and open-ended settings. Clues are fairly provided by default, * indicates w/o provided.
@@
Zero-Shot Reasoning LLM
DeepSeek-R1 deepseek 49.75 3.23 54.09
OpenAI-o3 o3 53.77 3.20 55.35
@@
Gemini-2.5-Pro gemini 52.84 3.25 53.57
Claude-3.7-Sonnet claude 53.31 3.27 52.83
GPT-4o gpt4 54.38 3.28 54.41
@@
Minigpt4-video Minigpt4-video 35.85 2.88 41.60
VideoLLaVA videollava 36.26 2.94 42.80
PLLaVA pllava 44.38 3.11 47.93
Qwen2.5-VL-7B qwen 46.95 3.13 48.72
NViLA nvila 48.42 2.95 46.08
VideoChat2 videochat2 50.27 3.15 52.24
Fine-tuning Open-Source MLLM
PLLaVA pllava 43.82 3.17 52.46
Qwen2.5-VL-7B qwen 51.26 3.26 54.32
VideoChat2 videochat2 51.90 3.27 53.79
Self-IRM 54.50 3.28 55.31
IRM 55.14 (+4.87) 3.29 56.02
@@
Figure 6: Accuracy and Efficiency Trade-off Results. Our model achieves both optimal performance and high efficiency.
Figure 7: Visualizations of multi-choice setting of I-VQA. Results indicate the powerful reasoning ability of IRM.
5.1 Implementation Details
Architectures and Training. We utilize VideoChat2 videochat2 as baseline with UMT-L umt as visual encoder and Mistral-7B mistral as LLM. We initialize IRM after vision-language alignment and connection to conduct instruction tuning. We trained IRM on 8 A6000 GPUs for about 3 hours. We use GPT-4o by default for candidate clue generation during training and IRM for inference, while the Self-IRM indicates using IRM for both training and inference. More details are in the Appendix.Sec. C.
Datasets. We conduct experiments on the I-VQA and the PSAV dataset adsdataset . The PSAV includes 1,002 examples for identifying the best strategy among 12 advertisement persuasion strategies.
Evaluation Metrics. For the open-ended setting of I-VQA, we adopt the “Score” and “$\text{Acc}_{\text{open}}$” metric following previous VideoQA works based on LLMs videollava ; qwen ; pllava ; stllm , where prediction is scored (from 0 to 5) and judged true or false by GPT-3.5. Besides, the high coherence between “Score” and human is proved in the Appendix.Sec. H, with Pearson correlation coefficient pearson reaching a high score of 0.93. We adopt option “Accuracy” for multi-choice setting. For PSAV, we follow the metric in adsdataset : “Accuracy” to evaluate the answer precision. Moreover, since multiple correct answers may exist for PSAV, we introduce “Recall” to evaluate the proportion of the answers covered.
Comparative Methods. We compare IRM with powerful Proprietary ImageLLMs GPT-4o gpt4o , Gemini-2.5-Pro gemini , and Claude-3.7-Sonnet claude . Similarly, we evaluate the leading open-source MLLMs, VideoLLaVA videollava , PLLaVA pllava , Minigpt4-video Minigpt4-video , NViLA nvila , Qwen2.5-VL qwen , and VideoChat2 videochat2 . We also evaluate reasoning models OpenAI-o3 o3 and DeepSeek-R1 deepseek . For fairer comparison, we fine-tune PLLaVA, Qwen2.5-VL, and VideoChat2. To ensure fairness, all comparison models are also prompted with the same context clues generated by GPT-4o during reasoning, which is strictly aligned with our IRM. Fine-tuning details and prompts for IRM and other models are in the Appendix.Sec. D, specific API version details are in the Appendix.Sec. C.
@@ 5.2 Main Results
Experiments of Invisible Evidence Questions.
As shown in Tab. 1, our IRM achieves SOTA in both the multi-choice and open-ended settings of the I-VQA dataset, with an accuracy of 55.14 and a GPT score of 3.29. This surpasses existing MLLMs, including fine-tuned VideoChat2 videochat2 , as well as GPT-4o gpt4 . Specifically, IRM demonstrates improvements of over 0.67% in $\text{Acc}_{\text{open}}$ and 0.76% in accuracy compared to the closest OpenAI-o3 and GPT-4o, exhibiting slightly stronger implicit reasoning abilities than strong commercial models. Accuracy and efficiency trade-off is shown in Fig. 5, detailed statistics are in the Appendix.Sec. B.
@@
Reasoning LLM OpenAI-o3 and DeepSeek-R1 demonstrate strong reasoning abilities under the enhancement of contextual clues generated by GPT-4o, successful and failed examples are in the Appendix.Sec. F. For validating the exclusion of commonsense questions, we also evaluate two reasoning models only with the question and options, the accuracy drops to 23.51 and 22.74, which is only slightly higher than random guessing. This result also indicates a small language bias perception of I-VQA. Visualizations of different settings in Fig. 7 and Fig. 5.2 demonstrate that IRM successfully infers the correct answer with dual refined clues. We also present the subjective results under the open-ended condition for examples that are not included in the open-ended subset. As shown in Fig. 5.2, IRM effectively generates multiple reasonable outcomes along with their probabilities.
Zero-shot Experiments of Implicit Intent Questions.
As shown in Tab. 2, IRM achieves SOTA performance with a recall of 45.41% and an accuracy of 46.08%, significantly outperforming the VideoChat2 by 24.49% and 6.86%, respectively. The results demonstrate IRM’s strong generalization ability and indicating we greatly enhance the reasoning capacity for general implicit questions. Visualizations are in Fig. 10 and the Appendix.Sec. A.
Figure 8: Visualization of the Open-ended Setting of I-VQA. Result indicates the reasoning and clues refining ability of IRM.
Figure 9: Open-ended Result of Example with Non-unique Answers. IRM effectively generates multi answers with probabilities.
Table 2: Results on PSAV Dataset. Our IRM also achieves SOTA performances in the zero-shot test.
Method Recall Accuracy
GPT-3.5 Story + Flan-t5 Classifier adsdataset - 33.41
GPT-3.5 Story + GPT-3.5 Classifier adsdataset - 35.02
OpenAI-o3 o3 36.38 45.84
GPT-4o gpt4 33.51 44.92
DeepSeek-R1 deepseek 23.62 21.96
VideoChat2 videochat2 20.92 39.22
PLLaVA pllava 38.36 39.40
NViLA nvila 38.13 41.09
Self-IRM 39.02 45.14
IRM 45.41 (+24.49) 46.08 (+6.86)
Figure 10: Visualizations of the PSAV Dataset. The results demonstrate the powerful implicit reasoning ability of IRM.

### v2

@@ 4 Methods
In this section, we present IRM as shown in Fig. 5, which mainly consists of the visual encoder, AIM (Action-Intent Module) (Sec. 4.1), VEM (Visual Enhanced Module) (Sec. 4.2), and the LLM. The AIM module is responsible for inferring clues relevant to the implicit question, while VEM enhances the visual representations of the clues inferred by AIM. Finally, the multimodal clue representations output by both modules are fed into the LLM to derive the answer to the final question.
Figure 5: IRM Framework. IRM employs dual-stream modeling of contextual actions and intent clues as implicit reasoning chains. Dual clues are generated and refined by AIM through hallucination mitigation and clue-question relationship deduction. Meanwhile, VEM takes the refined clues to enhance visual input. Finally, LLM is prompted with multi-modal dual clues to conduct implicit reasoning.
@@ 4.1 Action Intent Module
As discussed above, drawing on theories from cognitive science, we construct contextual clue information as an additional reasoning pathway to assist in implicit reasoning tasks. Furthermore, the visual evidence in the context is typically explicit. Consequently, extracting these direct, shallow intentions from actions is much less challenging than directly deducing the indirect, implicit intentions behind the question. Therefore, we proposed the AIM (Action Intent Module), which includes the extraction and refinement of Action Intent clues.
For the extraction of clues, we prompt our model (which can be substituted with other MLLMs, such as GPT-4o, during training) to leverage the contextual visual information to extract this explicit evidence and formulate it into candidate contextual clues structured as ”action-intent” pairs. Specifically, we prompt it with instruction prompt P (shown in the left-top legend of Fig. 5P: Task: Please extract …) to generate action-intent clue candidates.
LLM is prompted with the uniformly sampled frames of implicit visual input $v$ and the In-Context Learning prompt. Context action sets $\mathbb{A}_{c}=\{act^{0},...,act^{n}\}$ and corresponding intents set $\mathbb{I}=\{int^{0},...,int^{n}\}$ are generated, which are then embedded and concatenated as follows:
$\begin{split}X_{A}&=\texttt{Cat}(f_{T}(act^{0}),...,f_{T}(act^{n})),\\ X_{I}&=\texttt{Cat}(f_{T}(int^{0}),...,f_{T}(int^{n}))\end{split}$ (1)
where $f_{T}$ indicates the word embedding layer, Cat indicates concatenate.
To further refine the clues, we focus on mitigating the hallucination problem of clues while reasoning out key clues beneficial for implicit reasoning. The current action clues may be influenced by hallucinations, but this can be effectively mitigated through careful verification using visual information [MECD, hallucination, mmvu]. Therefore, we incorporate cross-action-vision attention to query each action candidate with global visual information $Q_{V}$ extracted by QFormer. Specifically, the mechanism is performed by the following equation:
$X_{I}^{V}=\text{$\Phi_{h}^{C}$}(X_{A},f_{P}(Q_{V}))$ (2)
@@
Moreover, action-intent clue pairs contribute differently to the reasoning of the implicit question. Therefore, we introduce the causal relations classifier with implicit question $q$ to determine which intent-action pairs assist in reasoning. Specifically, we concatenate the visually enhanced $X_{I}^{V}$ and the original intent tensor $X_{I}$ with the embedded question $f_{T}(q)$ and then pass them to the causal attention classifier [MECD] for relationship deduction. This process is represented as:
@@
$\mathcal{L}_{R}=\sum_{j=1}^{M}\mathcal{L}_{ce}(\hat{r}_{\hat{\sigma}(j)},r_{j})$ (4)
where $\mathcal{L}_{ce}$ is the cross-entropy loss. Since the set of $N$ predictions, $\{\hat{r}_{i}\}_{i=1}^{N}$, and the set of $M$ ground truth relation annotations, $\{r_{j}\}_{j=1}^{M}$, do not have a one-to-one correspondence, we first find an optimal bipartite matching between them. This matching, denoted as $\hat{\sigma}$, is found using an algorithm like the Hungarian algorithm to minimize the overall cost of pairing predictions to ground truths. The final loss is then the sum of the cross-entropy losses computed only over these optimally matched pairs, where $\hat{r}_{\hat{\sigma}(j)}$ is the prediction assigned to the ground truth $r_{j}$. During inference, there is no need to match against annotations; instead, the clue candidates are directly refined. The refined context action clues $X_{A}^{\prime}$ and intent clues $X_{I}^{\prime}$ can be represented as follows:
$\begin{split}X_{A}^{\prime}=\texttt{Cat}\left({X_{A}}[i]\right),X_{I}^{\prime}=\texttt{Cat}\left({X_{I}}[i]\right),\\ \hat{r}[i,0]>\hat{r}[i,1].\end{split}$ (5)
@@ 4.2 Visual Enhanced Module
As discussed above, the VEM first introduces QFormer [blip2] to compress extensive visual input $f_{V}(v)$ from the visual encoder, represented as $Q_{V}$. During the compression, the instruction to Qformer is: “Extract the context clues related to the question: q”. Following the previous works [videochat2, videollava, videollama, compress], compressed visual input $Q_{V}$ is then sent to visual projection layers $f_{P}$ to align with textual embeddings, represented by $X_{V}=f_{P}(Q_{V})$.
The aligned visual input is then sent to the AIM module for dual context clues reasoning. Meanwhile, the current visual extraction $X_{V}$ only implicitly suggests the need to extract key context visual clues. Therefore, for explicit visual enhancement, we establish a feedback path from AIM to VEM, introducing refined context clues generated by AIM to enhance the vanilla visual representation $X_{V}$ through a cross-attention mechanism:
$X_{V}^{\prime}=\text{$\Phi_{v}^{C}$}(X_{V},\texttt{Cat}(X_{A}^{\prime},X_{I}^{\prime})).$ (6)
@@
$Ans=\text{LLM}(X_{A}^{\prime},X_{I}^{\prime},X_{V}^{\prime},q).$ (7)
@@ 5 Experiments
5.1 Implementation Details
Architectures and Training. We utilize VideoChat2 [videochat2] as baseline with UMT-L [umt] as visual encoder and Mistral-7B [mistral] as LLM. We initialize IRM with the pre-trained VideoChat2 model after vision-language alignment, followed by an SFT stage to enhance the model’s implicit reasoning abilities. We trained IRM on 8 A6000 GPUs for about 3 hours. We use GPT-4o by default for candidate clue generation during the training phase of IRM. During inference, however, we use the same clue extraction prompt to have the IRM model generate the clues itself. More details are in the Appendix.Sec. C.
Datasets. We conduct experiments on the proposed I-VQA dataset and a representative advertising dataset, the PSAV dataset [adsdataset]. Advertising datasets are particularly well-suited for implicit reasoning, as they often employ subtle strategies to persuade the audience rather than explicitly instructing them to purchase a product or accept a premise. The PSAV dataset offers a curated collection of 1,002 video advertisements for analyzing persuasive tactics. Each video is annotated to identify the effective ones of 12 distinct advertising persuasion strategies.
Evaluation Metrics. For the evaluation of implicit video question answering, we follow the mainstream paradigms for video QA: open-ended and multiple-choice evaluation. For the open-ended setting of I-VQA, we adopt the “Score” and $\text{Acc}_{\text{open}}$” metric following previous VideoQA works based on LLMs [videollava, qwen, pllava, stllm], where prediction is scored (from 0 to 5) and judged true or false by GPT-3.5. Besides, the high coherence between “Score” and human is proved in the Appendix.Sec. G, with Pearson correlation coefficient [pearson] reaching a high score of 0.93. We adopt option “Accuracy” for multi-choice setting. For PSAV, we follow the metric in [adsdataset]: “Accuracy” to evaluate the answer precision. Moreover, since multiple correct answers may exist for PSAV, we introduce the complement metric “Recall” to evaluate the proportion of the answers covered.
Table 1: Results on I-VQA. IRM reaches SOTA performances in both multi-choice and open-ended settings. Clues are fairly provided by default; * indicates reasoning without clues provided. ZS indicates zero-shot, FT indicates fine-tuning.
@@ 5.1 Implementation Details
ZS Reasoning LLM
DeepSeek-R1 [deepseek] 49.75 3.23 54.09
[…]
@@
Gemini-2.5-Pro [gemini] 52.84 3.25 53.57
Claude-3.7-Sonnet [claude] 53.31 3.27 52.83
[…]
@@
Minigpt4-video [Minigpt4-video] 35.85 2.88 41.60
VideoLLaVA [videollava] 36.26 2.94 42.80
PLLaVA [pllava] 44.38 3.11 47.93
Qwen2.5-VL-7B [qwen] 46.95 3.13 48.72
NViLA [nvila] 48.42 2.95 46.08
[…]
FT Open-Source MLLM
PLLaVA [pllava] 43.82 3.17 52.46
Qwen2.5-VL-7B [qwen] 51.26 3.26 54.32
VideoChat2 [videochat2] 51.90 3.27 53.79
[…]
@@
Figure 6: Accuracy and Efficiency Trade-off Results. IRM achieves both optimal performance and high efficiency during reasoning.
Figure 7: Visualizations of multi-choice setting of I-VQA. The result indicates the powerful reasoning and clues refining ability of IRM.
Comparative Methods. We compare IRM with powerful Proprietary ImageLLMs GPT-4o [gpt4o], Gemini-2.5-Pro [gemini], and Claude-3.7-Sonnet [claude]. Similarly, we evaluate the leading open-source MLLMs, VideoLLaVA [videollava], PLLaVA [pllava], Minigpt4-video [Minigpt4-video], NViLA [nvila], Qwen2.5-VL [qwen], and VideoChat2 [videochat2]. We also evaluate reasoning models OpenAI-o3 [o3] and DeepSeek-R1 [deepseek]. For fairer comparison, we fine-tune the represented open-source MLLMs PLLaVA, Qwen2.5-VL, and VideoChat2 when conducting comparisons with our proposed IRM. To ensure fairness, all comparison models are also prompted with the same context clues generated by GPT-4o during reasoning, which is strictly aligned with our IRM. Fine-tuning details and prompts for IRM and other models are in the Appendix.Sec. D, specific API version details are in the Appendix.Sec. C.
@@ 5.2 Main Results
Evaluation of the I-VQA Dataset.
As shown in Tab. 1, our IRM achieves SOTA in both the multi-choice and open-ended settings of the I-VQA dataset, with an accuracy of 55.14 and a GPT score of 3.29. This surpasses existing MLLMs, including fine-tuned VideoChat2 [videochat2], as well as GPT-4o [gpt4]. Specifically, IRM demonstrates improvements of over 0.67% in $\text{Acc}_{\text{open}}$ and 0.76% in accuracy compared to the closest OpenAI-o3 and GPT-4o, exhibiting slightly stronger implicit reasoning abilities than strong commercial models. Accuracy and efficiency trade-off is shown in Fig. 5.1, and detailed statistics are in the Appendix.Sec. B. The results indicate that IRM achieves both optimal performance and high efficiency. Additionally, IRM with self-generated clues during both training and evaluation achieves the second-best performance, indicating that even under resource constraints and without relying on commercial MLLMs, the proposed clue-based strategies can effectively enhance the model’s implicit reasoning ability.
Figure 8: Visualization of the Open-ended Setting of I-VQA. GPT-4o linked removing the blindfold to starting hide-and-seek, missing the key clue that no one had disappeared, while IRM correctly reasoned the status.
Figure 9: Open-ended Result of Example with Non-unique Answers. In cases where hidden information has multiple possibilities, IRM effectively generates multiple answers with associated probabilities.
Table 2: Results on PSAV Dataset. IRM achieves SOTA performances in zero-shot test on unseen advertisement tasks.
Method Recall Accuracy
Story + Flan-t5 Classifier [adsdataset] - 33.41
Story + GPT-3.5 Classifier [adsdataset] - 35.02
@@
GPT-4o [gpt4] 33.51 44.92
DeepSeek-R1 [deepseek] 23.62 21.96
[…]
NViLA [nvila] 38.13 41.09
[…]
Figure 10: Visualizations of the PSAV Dataset. The results show that IRM accurately grasps the persuasion strategies.
The performances of $\text{GPT-4o}^{*}$ (GPT-4o without context clues provided, marked in red in Tab. 1) indicate the challenge of the implicit reasoning task. Besides, the fine-tuned MLLMs did not show a significant improvement over the zero-shot models, indicating that guiding the model’s reasoning process is far more crucial than simply exposing it to implicit questions. Conversely, our IRM significantly improves accuracy by 4.87% compared to the baseline, indicating the effectiveness of enhanced dual context clues.
Reasoning LLM OpenAI-o3 and DeepSeek-R1 demonstrate strong reasoning abilities under the enhancement of contextual clues generated by GPT-4o, successful and failed examples are in the Appendix.Sec. F. For validating the exclusion of commonsense questions, we also evaluate two reasoning models only with the question and options, the accuracy drops to 23.51 and 22.74, which is only slightly higher than random guessing. This result also indicates a minimal language bias [perception] of I-VQA. Visualizations of different settings in Fig. 7 and Fig. 5.2 demonstrate that IRM successfully infers the correct answer with dual refined clues. We also present the subjective results under the open-ended condition for examples that are not included in the open-ended subset. As shown in Fig. 5.2, IRM effectively generates multiple reasonable outcomes along with their probabilities.
Evaluation of the PSAV Dataset.
As detailed in Tab. 2, the IRM model has established a new state-of-the-art (SOTA) in the zero-shot generalization evaluation on the PSAV dataset. These results represent a significant improvement over VideoChat2, with IRM outperforming it by 24.49% in recall and 6.86% in accuracy. Although there are no video examples related to the advertising domain in the training set, we still achieved SOTA performance, demonstrating that our proposed method effectively enhances the reasoning capacity for general implicit questions. Visualizations are shown in Fig. 10 and the Appendix.Sec. A.